OpenStudy (anonymous):

If your starting salary were $40,000 and you received a 3% increase at the end of every year for 15 years, what would be the total amount, in dollars, you would have earned over the first 16 years that you worked? Round your answer to the nearest whole dollar.

OpenStudy (anonymous):

658000

OpenStudy (tkhunny):

Work it out. What's your plan? 40000 + 40000*(1.03) + 40000*(1.03)^2 + ... + 40000*(1.03)^15 Add it all up.

OpenStudy (anonymous):

where did you get 1.03?

OpenStudy (anonymous):

ok,so you do 40000x16,add 1200 on for every year for the next 15 years

OpenStudy (anonymous):

don't you use sn=a1(1-r^n)/1-r ?

OpenStudy (tkhunny):

"received a 3% increase " 40000*0.03 is the increase percentage each year. 40000*1.03 is the total after the increase each year.

OpenStudy (anonymous):

No if your looking for strategy then sorry,i just do these in my head and forget how to explain it elseware.

OpenStudy (tkhunny):

The formula will do you no good if you didn't see the value of 'r' in the problem statement. It will also do you no good if you're not sure it's the right one. Personally, and not everyone agrees with me on this, I refuse to learn those formulas. It is just as easy and just as fast to learn how to build them on the fly. Doing so absolutely eliminates these questions about which one and how to use it. Again, that is my opinion.

OpenStudy (tkhunny):

We've established that r = 1.03. What is n?

OpenStudy (anonymous):

16?

OpenStudy (tkhunny):

Let's see what you get.

OpenStudy (anonymous):

24187

OpenStudy (tkhunny):

What is that? It's at least $40,000 * 16 = $640,000 Please show your intermediate results. What is r^16?

OpenStudy (anonymous):

So i was right,why bother with the formula,(don't get me wrong,it great to learn) but i solved this problem with just common sense.

OpenStudy (tkhunny):

?? $658,000 is very short. Better get more common sense. :-)

OpenStudy (anonymous):

?,Ok,take a look,40000x16+18000=658,000

OpenStudy (anonymous):

18,000 is three percent over the next 15 years.

OpenStudy (anonymous):

s16=40,000(1-1.03^16) / 1-1.03=40,000(1-1.60) / -0.03=40000(-0.6) / -0.03 = -24000 / -0.03 = 800000 I failed.

OpenStudy (anonymous):

Hold that thought

OpenStudy (tkhunny):

Use more decimal places. That's all. 1.03^16 = 1.6047064390987875179380544409792 \(log_{10}(40000)=4.6ish\) so use at least 6 decimal places.

OpenStudy (anonymous):

24187?

OpenStudy (anonymous):

K,you'll have to explain how you keep getting such a low number.

OpenStudy (tkhunny):

You just said 800000. Then you reported the answer as 24187. That makes no sense at all. Do what you already did. Just use more decimal places. s16=40,000(1-1.03^16) / (1-1.03) Actually, with accumulation, it's a little easier to turn them around. Same answer, just less arithmetic to go wrong. s16=40,000(1.03^16 - 1) / (1.03 - 1)

OpenStudy (anonymous):

s16=40000(1-1.03^16) / 1- 1.03 = 40000(1-1.604706439) / -0.03 = 40000(-0.604706439) / -0.03 = -24188.25756 / -0.03 =806275 I got something different this time...

OpenStudy (tkhunny):

That's it. You will almost always get something different when you switch from making mistakes to NOT making mistakes.

OpenStudy (anonymous):

haha thanks :)

OpenStudy (tkhunny):

@deadmanwalking 40000x16+18000=658,000 This is strange. What is the 18000?

OpenStudy (anonymous):

me forgetting that i suck at percents,sorry,i also made a mistake and forgot to add a few numerals into certain parts of the equation.

OpenStudy (anonymous):

please ignore my posts completely if you haven't already.